OpenStudy (anonymous):

Chase scored 11 points on Monday, and he doubled his score each day thereafter. How many points did he score on Friday?

OpenStudy (anonymous):

I want to say 110 because if he has 11 on monday and you times that by 2 and times 5 you'll get 110

OpenStudy (anonymous):

monday=11 each incresing day doubled the points...so on tuesday=11+11=22 wednesday=22+22=44 thursday=44+44=88 friday=88+88=176

OpenStudy (anonymous):

get it..?